# expected information gain SLAM simultaneous localization and mapping

### v1

@@ VIII Discussion
With that said, this success can still be expanded to more fully demonstrate the model-building capability of tactile sensing. For example, while the current approach requires a convex object and end-effector to define a unique collision point, it should be possible to model non-convex objects as a union of convex shapes, further extending the space of possible geometry representations. Additionally, this work can be combined with a vision-based model-building solution [14, 18] to create a robust multi-modal system. This is a step towards the goal of a robot that can use all of the sensors at its disposal to model previously-unseen arbitrary objects in its environment throughout its entire deployment lifetime.
@@ Acknowledgments
This work was supported by an NSF CAREER Award under Grant No. FRR-2238480 and the RAI Institute. The authors extend their gratitude to Hrishikesh Sathyanarayan and Ian Abraham from the Intelligent Autonomy Lab at Yale University for all the insightful discussions early in the conception of this work.

### v2

@@ VIII Discussion
With that said, this success can still be expanded to more fully demonstrate the model-building capability of tactile sensing. For example, while the current approach requires a convex object and end-effector to define a unique collision point, it should be possible to model non-convex objects as a union of convex shapes, further extending the space of possible geometry representations. Additionally, this work can be combined with a vision-based model-building solution [jiang_fisherrf_2024, bianchini_vysics_2025] to create a robust multi-modal system. This is a step towards the goal of a robot that can use all of the sensors at its disposal to model previously-unseen arbitrary objects in its environment throughout its entire deployment lifetime.
@@ Acknowledgments
This work was supported by an NSF CAREER Award under Grant No. FRR-2238480 and the RAI Institute. The authors extend their gratitude to Hrishikesh Sathyanarayan and Ian Abraham from the CoCuRo Lab at the University of Sydney and Yale University for all the insightful discussions early in the conception of this work.